After absorbing 1.75 kJ of heat, a 0.454 kg block of nickel has a temperature of 36.3°C. What was its temperature before heating? (The specific heat of nickel is 0.440 J/gK

Give answer in Celsius
Answer #1

Given:

Q = 1.75 KJ = 1750 J

m = 0.454 Kg = 454 g

C = 0.44 J/g.oC

Tf = 36.3 oC

use:

Q = m*C*(Tf-Ti)

1750.0 = 454.0*0.44*(36.3-Ti)

36.3 - Ti = 8.8 oC

Ti = 27.5 oC

Answer: 27.5 oC